OpenStudy (anonymous):

Farmer Bob has pigs and chickens. He has 37 animals, and there are 124 legs among them altogether. How many chickens does Bob have?

OpenStudy (bibby):

piggies have 4 legs chickies have 2 legs 37 = 4p + 2c 124 = p + c

OpenStudy (anonymous):

do i use elimination with both of those equations?

OpenStudy (bibby):

You could. I'm using substitution right now

OpenStudy (anonymous):

what are you getting?

OpenStudy (bibby):

I'm not sure this is right. I'm getting 459/2 for c. I'm probably doing something wrong. Can you post your work for elimination?

OpenStudy (bibby):

Hopefully our new friends can help

OpenStudy (cwrw238):

one equation should be p + c = 37

OpenStudy (bibby):

rofl god dammit

OpenStudy (bibby):

looks like I mixed up my equations.

OpenStudy (cwrw238):

lol you got the legs mixed up with the animals

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

haha thank you!

OpenStudy (cwrw238):

2c + 2p = 74 2c + 4p = 124 subtracting 2p = 50 etc….

OpenStudy (bibby):

anyhow, 124 = 4p + 2c 37 = p + c c = 37 - p 2c = 74 - 2p 124 = 4p + 2c 124 = 2p + 74 50 = 2p p = 25

OpenStudy (anonymous):

so there are 12 chickens and 25 pigs. got it. Thanks for the help!

OpenStudy (the_fizicx99):

Lolll I hate these >.< I do them with random numbers

OpenStudy (bibby):

That's stupid lol. They give you the information you need

OpenStudy (the_fizicx99):

._. Nope, I have my method, much easier what you just did

OpenStudy (the_fizicx99):

Ofc, within the 37 animal limit